# AI-generated faces influence gender stereotypes and racial homogenization

### v2

@@ Results
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: User study results. Subfigures a to d summarize the participants’ responses in Studies 1 to 4, respectively. Boxes extend from the lower to upper quartile values, with a horizontal line at the median; whiskers extend to the most extreme values no further than 1.5 times the interquartile range from the box. P values are calculated using the t-test, unless one of the groups does not pass the Shapiro–Wilk test, in which case P values are calculated using the Mann-Whitney U test. p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001; ∗∗∗∗p$<$0.0001; $ns=$ not significant).
Additionally, for each question $Q_{i}:i\in\{1,2,3,4\}$, we recruited 135 participants from Prolific to answer $Q_{i}$ without being presented with the six AI-generated images. Figures 5a to 5d summarize the responses to questions $Q1$ to $Q4$, respectively. For all four questions, exposure to images generated by SDXL-Inc reduces bias compared to the baseline in which participants are not exposed to AI-generated images. In contrast, exposure to SDXL-generated images increases bias (compared to the baseline) for all questions apart from $Q_{1}$ where the effect is not significant. Next, we examine the effect of the AI label. As can be seen in the figure, there is no significant difference in participants’ responses when the images are labelled as being produced by an artist rather than being AI-generated. This holds for both SDXL and SDXL-Inclusive. Thus, we find no evidence that the AI label plays a role in the aforementioned phenomenon.
@@ Discussion
We set out to examine the stereotypes and biases in Stable Diffusion XL (SDXL), a text-to-image generator used daily by millions worldwide [8]. To the end, we developed a classifier to predict the race and gender of any given face image, and demonstrated that it achieves state-of-the-art performance. Using this classifier, we showed that the vast majority of faces generated by SDXL are White males (as previously shown by Ghosh and Caliskan [11]), while certain racial groups are extremely underrepresented, e.g., only 3% are Asian and only 5% are Indian. Biases were also found when considering various attributes, e.g., associating beauty with femininity and intelligence with masculinity. Some of these biases are less severe in the dataset on which Stable Diffusion was trained, suggesting that certain biases are further exacerbated by the model itself.
@@
Biased text-to-image models may contribute to the normalization of gender stereotypes, potentially shaping societal attitudes towards the roles and capabilities of women in various professions. For instance, as we have demonstrated, Stable Diffusion mostly depicts secretaries and nurses as women while depicting managers, doctors, and professors as men; similar gender stereotypes were reported in the literature [13, 10]. Bearing in mind that millions of people worldwide are already using such models daily, addressing gender stereotypes in these models can be crucial. As suggested by Study 2, such stereotypes can be reduced by an inclusive model and can be exacerbated by a non-inclusive one, indicated the potential of AI in alleviating gender inequality.
@@
Biased representations of gender and race may contribute to the creation of content that not only misrepresents certain groups but may also perpetuate discriminatory practices. This could be detrimental in the context of advertising, marketing, and media campaigns, where visuals hold substantial influence. For example, as our analysis has shown, Stable Diffusion associates low-income jobs such as Cleaner, Janitor, and Security Guard with Black people, while associating higher-prestige jobs such as Doctor, Lawyer, and Professor with White people. These findings reflect what has been reported by Bianchi et al. [9], showing that several of the most prestigious, high-paying professions are represented by Stable Diffusion (v1-4) as White. Another example of bias is the association between crime and Black people, as well as the assassination between terrorism and Middle Easterners, which may reinforce existing biases against these racial groups [23, 24]. As we have seen in Study 1, whether or not the model is inclusive can affect people’s perception of the racial distribution of certain professions, and this effect is likely to grow more pronounced as the use of AI-generated images becomes more widespread.
@@
Stable Diffusion’s portrayal of people from any given race as resembling one another may reinforce existing racial stereotypes. For instance, as we have demonstrated, Stable Diffusion mostly depicts Middle Eastern men as dark-skinned, bearded, and wearing a traditional headdress, and mostly depicts Middle Eastern women as dark-skinned, wearing a headscarf. Such oversimplified and generalized depictions of a particular racial group can be culturally insensitive, and may misrepresent the true diversity within that group. They may also lead to feelings of alienation, low self-esteem, and a sense of being misunderstood. As we have demonstrated in Studies 3 and 4, racial homogenization can be reduced using inclusive models, and can be exacerbated using non-inclusive ones.
@@ Methods
Similarly, we fine-tuned SDXL using LORA in the process of creating SDXL-Div—our solution for mitigating the lack of diverse facial features in images generated by SDXL. To this end, we created a dataset consisting of prompt-image pairs. Each such pair consisted of an image taken from the “Flickr-Faces-HQ dataset” (described above) along with the prompt: “a photo of $X$ person, looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”, where $X$ is the race depicted in the image. This process yielded the SDXL-Div model.
@@
I) Face detection: This stage is carried out using a Multi-task Cascaded Convolutional Neural Network (MTCNN) [30]. Specifically, MTCNN is a deep cascaded multitask framework that utilizes the inherent correlation between detection and alignment to improve the performance [30]. It leverages a cascaded architecture with three stages of deep convolutional networks to predict faces. MTCNN outperformed other methods across several challenging benchmarks. We selected this face detector in our work due to its ability to balance high detection accuracy and run-time speed. We configured the detector to exclude boundary-boxes with confidence scores $\leq 0.9$.
@@
III) Classification stage: This stage is carried out using a Support Vector Machine (SVM) [17]. More specifically, we trained two SVM classifiers to predict the race and gender using the embedding vectors extracted from the FairFace images during the previous stage. The hyper-parameters used in SVM are as follows: Regularization parameter C = 1; Kernel type = Radial Basis Function. This stage is repeated twice; once for race, and once for gender. However, the previous two stages are executed only once.
@@
N.A., T.R. and Y.Z. conceived the study, designed the research, produced the visualizations, and wrote the manuscript; N.A. performed the literature review, designed and implemented the classifiers (race and gender) and debiasing solutions (SDXL-Inc and SDXL-Div), collected and analyzed the data, and ran the experiments. T.R. and Y.Z. designed and ran the user study, and analyzed the study results.

### v3

@@ Results
Additionally, for each question $Q_{i}:i\in\{1,2,3,4\}$, we recruited 135 participants from Prolific to answer $Q_{i}$ without being presented with the six AI-generated images. Figures 5a to 5d summarize the responses to questions $Q1$ to $Q4$, respectively. For all four questions, exposure to images generated by SDXL-Inc reduces bias compared to the baseline in which participants are not exposed to AI-generated images. In contrast, exposure to SDXL-generated images increases bias (compared to the baseline) for all questions apart from $Q_{1}$ where the effect is not significant. Next, we examine the effect of the AI label. As can be seen in the figure, there is no significant difference in participants’ responses when the images are labeled as being produced by an artist rather than being AI-generated. This holds for both SDXL and SDXL-Inclusive. Thus, we find no evidence that the AI label plays a role in the aforementioned phenomenon, indicating that the observed effect is entirely driven by the image’s content, irrespective of its source.
Figure 5: Survey experiment results. Subfigures a to d summarize the participants’ responses in Studies 1 to 4, respectively. Boxes extend from the lower to upper quartile values, with a horizontal line at the median; whiskers extend to the most extreme values no further than 1.5 times the interquartile range from the box. P values are calculated using the t-test, unless one of the groups does not pass the Shapiro–Wilk test, in which case P values are calculated using the Mann-Whitney U test. p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001; ∗∗∗∗p$<$0.0001; $ns=$ not significant).
@@ Discussion
We set out to examine the stereotypes and biases in Stable Diffusion XL (SDXL), a text-to-image generator used daily by millions worldwide [8]. To this end, we developed a classifier to predict the race and gender of any given face image, and demonstrated that it achieves state-of-the-art performance. Using this classifier, we showed that the vast majority of faces generated by SDXL are White males (as previously shown by Ghosh and Caliskan [11]), while certain racial groups are rarely generated, e.g., only 3% are Asian and only 5% are Indian. Biases were also found when considering various attributes, e.g., associating beauty with femininity and intelligence with masculinity. Some of these biases are less severe in the dataset on which Stable Diffusion was trained, suggesting that certain biases are further exacerbated by the model itself.
@@
Biased text-to-image models may contribute to the normalization of gender stereotypes, potentially shaping societal attitudes towards the roles and capabilities of women in various professions. For instance, as we have demonstrated, Stable Diffusion mostly depicts secretaries and nurses as women while depicting managers, doctors, and professors as men; similar gender stereotypes were reported in the literature [13, 10]. Bearing in mind that millions of people worldwide are already using such models daily, addressing gender stereotypes in these models can be crucial. As suggested by Study 2, such stereotypes can be reduced by an inclusive model and can be exacerbated by a non-inclusive one, indicating the potential of AI in alleviating gender inequality.
@@
Biased representations of gender and race may contribute to the creation of content that not only misrepresents certain groups but may also perpetuate discriminatory practices. This could be detrimental in the context of advertising, marketing, and media campaigns, where visuals hold substantial influence. For example, as our analysis has shown, Stable Diffusion associates low-income jobs such as Cleaner, Janitor, and Security Guard with Black people, while associating higher-prestige jobs such as Doctor, Lawyer, and Professor with White people. These findings reflect what has been reported by Bianchi et al. [9], who showed that several of the most prestigious, high-paying professions are represented by Stable Diffusion (v1-4) as White. This is consistent with the historical and ongoing processes of occupational segregation based on race and gender, which could be linked to the concept of stereotype threat [44], where awareness of negative stereotypes can negatively impact the performance of individuals from stereotyped groups [16]. Another example of bias is the association between crime and Black people, as well as the assassination between terrorism and Middle Easterners, which may reinforce existing biases against these racial groups [40, 41]. As we have seen in Study 1, whether or not the model is inclusive can affect people’s perception of the racial distribution of certain professions, and this effect is likely to grow more pronounced as the use of AI-generated images becomes more widespread.
@@
Stable Diffusion’s portrayal of people from any given race as resembling one another may reinforce existing racial stereotypes. For instance, as we have demonstrated, Stable Diffusion predominently depicts Middle Eastern men as bearded and brown-skinned, while depicting Middle Eastern women as mostly brown-skinned and wearing traditional attire. This echoes the concept of Orientalism [18], which critiques the simplified portrayals of the East, particularly the Middle East, in Western representations. Such oversimplified and generalized depictions of a particular racial group can be culturally insensitive, and may misrepresent the true diversity within that group. They may also lead to feelings of alienation, low self-esteem, and a sense of being misunderstood. This can be understood in the context of social comparison theory [19], where individuals compare themselves to others to evaluate their own abilities and opinions. When exposed to limited and stereotypical representations, it can negatively affect their self-esteem and sense of belonging. As we have demonstrated in Studies 3 and 4, racial homogenization can be reduced using inclusive models, and can be exacerbated using non-inclusive ones.
@@ Methods
Similarly, we fine-tuned SDXL using LORA in the process of creating SDXL-Div—our solution for mitigating racial and gender homogenization in SDXL. To this end, we created a dataset consisting of prompt-image pairs. Each such pair consisted of an image taken from the “Flickr-Faces-HQ dataset” (described above) along with the prompt: “a photo of $X$ person, looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”, where $X$ is the race depicted in the image. This process yielded the SDXL-Div model.
@@
I) Face detection: This stage is carried out using a Multi-task Cascaded Convolutional Neural Network (MTCNN), which is a deep cascaded multitask framework that utilizes the inherent correlation between detection and alignment to improve performance [48]. It leverages a cascaded architecture with three stages of deep convolutional networks to predict faces. MTCNN outperformed other methods across several challenging benchmarks. We selected this face detector in our work due to its ability to balance high detection accuracy and run-time speed. We configured the detector to exclude boundary-boxes with confidence scores $\leq 0.9$.
@@
III) Classification: This stage is carried out using a Support Vector Machine (SVM) [24]. More specifically, we trained two SVM classifiers to predict the race and gender using the embedding vectors extracted from the FairFace images during the previous stage. The hyper-parameters used in SVM are as follows: Regularization parameter C = 1; Kernel type = Radial Basis Function. This stage is repeated twice; once for race, and once for gender. However, the previous two stages are executed only once.
@@
Data Availability
All of the data used in our analysis can be found at the following repository: https://github.com/comnetsAD/AI-generated-faces.
@@
N.A., T.R. and Y.Z. conceived the study, designed the research, produced the visualizations, and wrote the manuscript; N.A. performed the literature review, designed and implemented the classifiers (race and gender) and debiasing solutions (SDXL-Inc and SDXL-Div), collected and analyzed the data, and ran the experiments. T.R. and Y.Z. designed and ran the survey experiment and analyzed the results.